# Dust evolution during the protostellar collapse: influence on the coupling between the neutral gas and the magnetic field

### v1

@@ 5.2 A hint at a magnetic flux regulation
The ambipolar drift can either be enhanced by increasing the magnetic field strength or $\delta_{\mathrm{AD}}$. When fragmentation is inefficient ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$), we can see on Fig. 6 that in increasing the ambipolar drift intensity, the small grains are more swiftly depleted and thus the ambipolar diffusion timescale drops more rapidly. In addition, Fig. 5 shows that a higher magnetic field allows the ambipolar resistivity $\eta_{\mathrm{AD}}$ to rise more rapidly, which entails an ambipolar diffusion timescale dropping faster. Therefore we can identify here a mechanism of magnetic flux regulation. The higher the magnetic field, the stronger the ambipolar drift, the faster the small grains are removed and thus the faster the magnetic field diffuses. However when considering an efficient fragmentation ($\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, dotted lines), the trend is the same except for high values of the ambipolar diffusion intensity. Indeed, in going from $\delta_{\mathrm{AD}}=1$ to $\delta_{\mathrm{AD}}=10$, it takes more time for the ambipolar diffusion timescale to drop, thus suppressing the magnetic flux regulation mechanism. This is due to the fact that for such fragile grains and high $\delta_{\mathrm{AD}}$, the ambipolar diffusion drift between dust grains induces destructive collisions leading to a replenishment of the small grains instead of a depletion of the latter.
@@ 5.3 Saturation of the magnetic field
Finally, Fig. 7 displays the evolution of the ambipolar diffusion timescale for different values of the initial magnetic field strength and ionization rate induced by cosmic rays. The black line represents the free-fall timescale. The ambipolar diffusion timescale is computed via Eq. (22) where each gas density is associated to a radius $r$ assuming a Singular Isothermal Sphere (see Shu, 1977)
@@
We can see that for a given value of the ionization rate $\zeta$, there exists a critical value of the initial magnetic field for which the ambipolar diffusion timescale will inexorably drop to equate the freefall timescale. When it does so, the magnetic field diffuses in such a way as to remain constant. The larger the cosmic-rays ionization rate, the better is the coupling between the plasma and the magnetic field (i.e., the lower the ambipolar resistivity $\eta_{\mathrm{AD}}$). Consequently a higher initial magnetic field $B_{\mathrm{0}}$ is needed to produce an ambipolar drift strong enough to remove the small grains and thus to allow the magnetic field to diffuse significantly (i.e., to ensure that $t_{\mathrm{AD}}=t_{\mathrm{ff}}$).
In such circumstances, we can define a saturation magnetic field that cannot be exceeded. We start from $t_{\mathrm{AD}}=t_{\mathrm{ff}}$ and assume that the ambipolar resistivity $\eta_{\mathrm{AD}}$ is controlled by the ion contribution, just as we did before. Using the analytical expression for $\eta_{\mathrm{AD}}$ which is plotted on Fig. 5, the previously defined expression for the ion reduced temperature, the free-fall timescale defined in Eq. (7) and Eq. (25), we get
@@
Whatever the initial magnetic field strength $B_{\mathrm{0}}$, the small grains will be removed, preventing the magnetic field strength from exceeding the value deduced from Eq. (28). We stress that this expression is valid as long as the dust contribution is absent, that is to say prior to a potential significant replenishment of the small grains by fragmentation, and in the first instants of the collapse once they have been removed. Note however that at high densities in protoplanetary disks, the fragmentation barrier could be potentially reached after a few orbital periods, allowing the small grain populations to be replenished and consequently allowing the gas and the magnetic field to decouple once more (in regions where the temperature is below the sublimation temperature of the dust ($T\sim 1500\ \mathrm{K}$) and where the ionization induced by the protostar is negligible). Also, with rotation and magnetic pressure, the free-fall timescale should be longer and therefore diffusion of the magnetic flux would be even easier.
@@ 6 Discussion
6.1 Is there a small grain depletion catastrophe?
In light of the soaring of the ambipolar resistivity $\eta_{\mathrm{AD}}$ at low density resulting from dust coagulation, one can wonder if small grain depletion has catastrophic consequences on protoplanetary disk formation?
6.1.1 Tensions with observed disks size
@@ 6.1.2 Ways to preserve the small grain population
To alleviate the tension with observations, we identified several mechanisms preventing the small grains from being too severely depleted, thus allowing to reduce the ambipolar resistivity $\eta_{\mathrm{AD}}$ prior to $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$:
@@
  • Low grain-grain sticking efficiency: Likewise the surface energy of the grains, this parameter remains poorly constrained. As depicted on Fig. 17, allowing for one collision event out of ten to result in grain sticking permits to recover an MRN-like ambipolar resistivity $\eta_{\mathrm{AD}}$ profile. In this case, the depletion of small grains is substantially curtailed.
@@
  • Electrostatic repulsion: With the reference value $\delta_{\mathrm{AD}}=1$, the ambipolar drift generates grain collisional velocities larger than the velocity threshold imposed by the Coulomb barrier (see Eq. 18). However, this is no longer the case when reducing the ambipolar drift intensity to $\delta_{\mathrm{AD}}=0.1$. As a consequence, small grain depletion is further hindered and the combined effect of a lower ambipolar drift intensity and grain electrostatic repulsion leads to a significant reduction of the ambipolar resistivity $\eta_{\mathrm{AD}}$ (see Fig. 4). Similarly, keeping $\delta_{\mathrm{AD}}=1$, we notice that reducing the initial magnetic field generates a weaker ambipolar drift. Our simulations show that for an initial magnetic field strength (at $n_{\mathrm{H}}=10^{4}\ \mathrm{cm^{-3}}$) of $B_{0}\leq B_{\mathrm{0,min}}=10\ \mathrm{\mu G}$ (for $\zeta=5\times 10^{-17}\ \mathrm{s^{-1}}$ and $\delta_{\mathrm{AD}}=1$, that is for an initial current strength of $\delta_{\mathrm{AD}}\frac{|\mathbf{B_{\mathrm{0,min}}}|}{\lambda_{\mathrm{J}}}=5.72\times 10^{-4}\ \mathrm{\mu G\ AU^{-1}}$), electrostatic repulsion effectively mitigates small grain coagulation and leads to a substantially lower ambipolar resistivity $\eta_{\mathrm{AD}}$. Equation (21) is then a good approximation and the magnetic field can amplify in the early phase of the collapse. The value $B_{\mathrm{0,min}}$ below which electrostatic repulsion comes into play is somehow almost insensitive to the cosmic ionization rate. We measured values between $9.5\ \mathrm{\mu G}$ and $13.4\ \mathrm{\mu G}$ for $\zeta$ in the range $\left[5\times 10^{-18},5\times 10^{-16}\right]\ \mathrm{s^{-1}}$. However, it should be kept in mind that although the population of small grains is preserved and the ambipolar resistivity is significantly reduced, a lower initial magnetic field (i.e., a higher mass-to-flux ratio) entails a weaker magnetic braking.
@@
Relying on electrostatic repulsion or erosion, we see from Fig. 4 that in either scenario, the difference between the computed ambipolar resistivity value and that associated to a MRN dust distribution is of one order of magnitude at most (the ratio ranges between 2 at low density and 10 shortly before $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$). Relying on the analytical formulae for the radius of the early protoplanetary disk derived in Hennebelle et al. (2016) ($r_{\mathrm{AD}}\propto\eta_{\mathrm{AD}}^{\frac{2}{9}}$, and see Lee et al. (2024) for an extension to misalignment between magnetic field and rotation axis), we find that the corresponding difference in disk size would be of a factor of $10^{\frac{2}{9}}\sim 1.67$ in the worst case, in contrast to a factor of $100^{\frac{2}{9}}\sim 2.78$ without repulsion or erosion. Comparing now the MRN case with the coagulation-repulsion case with a lower sticking efficiency ($\epsilon_{\mathrm{stick}}=0.5$), we find a factor of $2^{\frac{2}{9}}\sim 1.16$. In terms of actual disk size values, we gather in Table 2 the comparison between the different scenarios and readily see that we obtain reasonable disk radii with either repulsion or erosion. The initial vertical magnetic field is assumed to be $B_{\mathrm{z}}=30\ \mathrm{\mu G}$, and the star/disk system mass to be $M=0.1\ M_{\odot}$ (consistent with a class 0 disk produced by the collapse of a one solar mass dense core). The ambipolar resistivity values $\eta_{\mathrm{AD}}$ are taken at the point of maximum departure from the MRN value, i.e., slightly before $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$.
Disk radius
@@
Table 2: Disk radius $r_{\mathrm{AD}}$ (Hennebelle et al., 2016) for different scenarios, with a fixed MRN dust distribution, accounting for coagulation only, coagulation and electrostatic repulsion, coagulation and erosion, coagulation-repulsion and a lower sticking efficiency ($\epsilon_{\mathrm{stick}}=0.5$). The initial vertical magnetic field is assumed to be $B_{\mathrm{z}}=30\ \mathrm{\mu G}$, and the star-disk system mass to be $M=0.1\ M_{\odot}$. The ambipolar resistivity values $\eta_{\mathrm{AD}}$ are taken at the point of maximum departure from the MRN value, i.e., slightly before $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$.
@@
6.2 Resilient grains or fragile grains? Uncertainties in the collision outcomes
@@
Uncertainties remain in regard to the description of collision outcomes. We recall here that some of them have been ignored in this work (see Wurm & Teiser (2021); Blum (2018) for a detailed list). Bouncing for example, is expected to occur prior to fragmentation, and leads to compaction in place of grain sticking. The bouncing barrier has been widely described in the literature (Zsom et al., 2010). Erosion as a consequence of grain-gas drag could replenish the system with small grains but is expected to be relevant mainly in protoplanetary disks where dust grains can experience a strong gas head wind (Rozner et al., 2020). More relevant to our environment is grain-grain erosion (Gundlach & Blum, 2015; Schrapler et al., 2018) whose influence has been investigated in this work. We based our model on the experimental results of Schrapler et al. (2018). In this regard, we found that grain-grain erosion is unconditionally very efficient at high density, where collisional velocities and size ratios are large, leading to noticeable variations in the Ohm and ambipolar resistivities. However, it is efficient at low density only for sufficiently strong ambipolar drifts. Therefore, given the essential role of this process, future laboratory experiments will have to reinforce our understanding of this collision outcome and to further constrain the velocity threshold and efficiency. Additionally, electrostatic and magnetic reactions resulting from the charges carried by the grains may affect dust growth. In UV-shielded regions (grain photoionization negligible) where the electrons are abundant ($n_{\mathrm{e}}>n_{\mathrm{dust}}$, i.e., efficient gas ionization), plasma charging dominates and dust grains are expected to be negatively charged on average (Draine & Sutin, 1987), electrostatic forces thus being repulsive and detrimental to grain coagulation (Okuzumi, 2009; Akimkin et al., 2020). When accounting for dust fragmentation and electrostatic repulsion, the population of sub-micron grains was found to be enhanced and the dust coagulation/fragmentation equilibrium to be severely affected (see Akimkin et al. (2023) who consider hot UV-shielded regions of protoplanetary disks where gas thermal ionization is important, with $T=1000\ \mathrm{K}$ and $\rho=10^{-12}\ \mathrm{g\ cm^{-3}}$). In the context of protostellar collapse, since the densities involved are very low, we believe that between two consecutive grain-grain collisions, the dust particles would have sufficient time to settle back to their average equilibrium negative charge value imposed by their interaction with electrons and ions. In this case, dust grains would feel repulsive electrostatic forces. As discussed previously, this can curtail to some extent the very efficient coagulation observed at the very beginning of the collapse.
@@
Moreover, the conditions needed for sticking and fragmentation are diverse. Those relevant to a dense core environment are not well constrained in the literature either, and not well reproduced in laboratory experiments. That is why we explored a wide range of grain surface energies $\gamma_{\mathrm{grain}}$. For instance, collision outcomes for icy-grains have turned out to be temperature dependent (Gärtner et al., 2017). Also, icy-grains have been believed to be more sticky that bare-silicates grains (Gundlach et al., 2011; Drazkowska, 2017), thus enhancing coagulation and being less prone to bouncing and compaction. Nevertheless, recent studies revealed that the grain surface energy $\gamma_{\mathrm{grain}}$ (which controls both the sticking efficiency and the fragmentation threshold) of icy-grains could sharply decrease from $T=200\ \mathrm{K}$ to $T=175\ \mathrm{K}$ due to a change in the crystal structure (again within the frame of the laboratory experiments, where $P=1\ \mathrm{bar}$ and consequently the ice melting temperature being $T=273\ \mathrm{K}$). Their measurements suggested a grain surface energy ranging between $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$ at low temperature and $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$ at higher temperature, which are the limit values we used in this work. For more details, refer to Musiolik & Wurm (2019). Note however that amorphous ice at low pressure and temperature could recover an efficient sticking under UV irradiation (Musiolik, 2021). We explored the impact of the grain sticking efficiency in Sect. 17 and found that values above $0.3$ do not prevent the population of small grains from being depleted at low densities, thus allowing the ambipolar resistivity $\eta_{\mathrm{AD}}$ to dwell on very large values.

### v2

@@ 5.2 A hint at a magnetic flux regulation
The ambipolar drift can be enhanced by either increasing the magnetic field strength or raising $\delta_{\mathrm{AD}}$. When fragmentation is inefficient ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$), we can see on Fig. 6 that in increasing the ambipolar drift intensity, the small grains are more swiftly depleted and thus the ambipolar diffusion timescale drops more rapidly. In addition, Fig. 5 shows that a higher magnetic field allows for the ambipolar resistivity, $\eta_{\mathrm{AD}}$, to rise more rapidly, which entails an ambipolar diffusion timescale dropping faster. Therefore we can identify here a mechanism of magnetic flux regulation. The higher the magnetic field, the stronger the ambipolar drift, the faster the small grains are removed and thus the faster the magnetic field diffuses. However when considering an efficient fragmentation ($\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, dotted lines), the trend is the same except for high values of the ambipolar diffusion intensity. Indeed, in going from $\delta_{\mathrm{AD}}=1$ to $\delta_{\mathrm{AD}}=10$, it takes more time for the ambipolar diffusion timescale to drop, thus suppressing the magnetic flux regulation mechanism. This is due to the fact that for such fragile grains and high $\delta_{\mathrm{AD}}$, the ambipolar diffusion drift between dust grains induces destructive collisions leading to a replenishment of the small grains instead of a depletion of the latter.
@@ 5.3 Saturation of the magnetic field
Finally, Fig. 7 displays the evolution of the ambipolar diffusion timescale for different values of the initial magnetic field strength and ionisation rate induced by cosmic rays. The black line represents the free-fall timescale. The ambipolar diffusion timescale is computed via Eq. (22) where each gas density is associated with a radius, $r,$ assuming a singular isothermal sphere (see Shu, 1977)
@@
We can see that for a given value of the ionisation rate $\zeta$, there exists a critical value of the initial magnetic field for which the ambipolar diffusion timescale will inexorably drop to equate the freefall timescale. When it does so, the magnetic field diffuses in such a way as to remain constant. The larger the cosmic-rays ionisation rate, the better is the coupling between the plasma and the magnetic field (i.e. the lower the ambipolar resistivity, $\eta_{\mathrm{AD}}$). Consequently a higher initial magnetic field $B_{\mathrm{0}}$ is needed to produce an ambipolar drift strong enough to remove the small grains and, thus, to allow the magnetic field to diffuse significantly (i.e. to ensure that $t_{\mathrm{AD}}=t_{\mathrm{ff}}$).
In such circumstances, we can define a saturation magnetic field that cannot be exceeded. We start from $t_{\mathrm{AD}}=t_{\mathrm{ff}}$ and assume that the ambipolar resistivity $\eta_{\mathrm{AD}}$ is controlled by the ion contribution, just as we did before. Using the analytical expression for $\eta_{\mathrm{AD}}$ which is plotted on Fig. 5, the previously defined expression for the ion reduced temperature, the free-fall timescale defined in Eq. (7) and Eq. (25), we get:
@@
Whatever the initial magnetic field strength $B_{\mathrm{0}}$, the small grains will be removed, preventing the magnetic field strength from exceeding the value deduced from Eq. (28). We stress that this expression is valid as long as the dust contribution is absent, that is to say prior to a potential significant replenishment of the small grains by fragmentation, and in the first instants of the collapse once they have been removed. We note, however, that at high densities in protoplanetary discs, the fragmentation barrier could be potentially reached after a few orbital periods, allowing for the small grain populations to be replenished and consequently allowing the gas and the magnetic field to decouple once more – in regions where the temperature is below the sublimation temperature of the dust ($T\sim 1500\ \mathrm{K}$) and where the ionisation induced by the protostar is negligible. Also, with rotation and magnetic pressure, the free-fall timescale should be longer and therefore diffusion of the magnetic flux would be even easier.
@@ 6 Discussion
6.1 Considering whether there is a small grain depletion catastrophe
In light of the soaring of the ambipolar resistivity, $\eta_{\mathrm{AD}}$, at low density resulting from dust coagulation, we may wonder whether small grain depletion has catastrophic consequences on protoplanetary disc formation.
6.1.1 Tensions with observed discs size
@@ 6.1.2 Ways to preserve the small grain population
To alleviate the tension with observations, we identified several mechanisms preventing the small grains from being too severely depleted, thereby allowing us to reduce the ambipolar resistivity $\eta_{\mathrm{AD}}$ prior to $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$.
@@
Low grain-grain sticking efficiency: Likewise the surface energy of the grains, this parameter remains poorly constrained. As depicted on Fig. 17, allowing for one collision event out of ten to result in grain sticking allows us to recover an MRN-like ambipolar resistivity $\eta_{\mathrm{AD}}$ profile. In this case, the depletion of small grains is substantially curtailed.
@@
Electrostatic repulsion: With the reference value $\delta_{\mathrm{AD}}=1$, the ambipolar drift generates grain collisional velocities larger than the velocity threshold imposed by the Coulomb barrier (see Eq. 18). However, this is no longer the case when reducing the ambipolar drift intensity to $\delta_{\mathrm{AD}}=0.1$. As a consequence, small grain depletion is further hindered and the combined effect of a lower ambipolar drift intensity and grain electrostatic repulsion leads to a significant reduction of the ambipolar resistivity $\eta_{\mathrm{AD}}$ (see Fig. 4). Similarly, keeping $\delta_{\mathrm{AD}}=1$, we notice that reducing the initial magnetic field generates a weaker ambipolar drift. Our simulations show that for an initial magnetic field strength (at $n_{\mathrm{H}}=10^{4}\ \mathrm{cm^{-3}}$) of $B_{0}\leq B_{\mathrm{0,min}}=10\ \mathrm{\mu G}$ (for $\zeta=5\times 10^{-17}\ \mathrm{s^{-1}}$ and $\delta_{\mathrm{AD}}=1$, that is for an initial current strength of $\delta_{\mathrm{AD}}\frac{|\mathbf{B_{\mathrm{0,min}}}|}{\lambda_{\mathrm{J}}}=5.72\times 10^{-4}\ \mathrm{\mu G\ AU^{-1}}$), electrostatic repulsion effectively mitigates small grain coagulation and leads to a substantially lower ambipolar resistivity $\eta_{\mathrm{AD}}$. Equation (21) is then a good approximation and the magnetic field can amplify in the early phase of the collapse. The value $B_{\mathrm{0,min}}$ below which electrostatic repulsion comes into play is somehow almost insensitive to the cosmic ionisation rate. We measured values between $9.5\ \mathrm{\mu G}$ and $13.4\ \mathrm{\mu G}$ for $\zeta$ in the range $\left[5\times 10^{-18},5\times 10^{-16}\right]\ \mathrm{s^{-1}}$. However, it should be kept in mind that although the population of small grains is preserved and the ambipolar resistivity is significantly reduced, a lower initial magnetic field (i.e. a higher mass-to-flux ratio) entails a weaker magnetic braking.
@@
Relying on electrostatic repulsion or erosion, we see from Fig. 4 that in either scenario, the difference between the computed ambipolar resistivity value and that associated with an MRN dust distribution is of one order of magnitude at most (the ratio ranges between 2 at low density and 10 shortly before $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$). Relying on the analytical formulae for the radius of the early protoplanetary disc derived in Hennebelle et al. (2016) ($r_{\mathrm{AD}}\propto\eta_{\mathrm{AD}}^{\frac{2}{9}}$, and see Lee et al. (2024) for an extension to misalignment between magnetic field and rotation axis), we find that the corresponding difference in disc size would be of a factor of $10^{\frac{2}{9}}\sim 1.67$ in the worst case, in contrast to a factor of $100^{\frac{2}{9}}\sim 2.78$ without repulsion or erosion. Comparing now the MRN case with the coagulation-repulsion case with a lower sticking efficiency ($\epsilon_{\mathrm{stick}}=0.5$), we find a factor of $2^{\frac{2}{9}}\sim 1.16$. In terms of actual disc size values, we gather in Table 2 the comparison between the different scenarios and readily see that we obtain reasonable disc radii with either repulsion or erosion. The initial vertical magnetic field is assumed to be $B_{\mathrm{z}}=30\ \mathrm{\mu G}$, and the star/disc system mass to be $M=0.1\ M_{\odot}$ (consistent with a class 0 disc produced by the collapse of a one solar mass dense core). The ambipolar resistivity values $\eta_{\mathrm{AD}}$ are taken at the point of maximum departure from the MRN value, namely, slightly before $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$.
Table 2: disc radius $r_{\mathrm{AD}}$ for different scenarios.
disc radius
@@
 [^2] Disc radius $r_{\mathrm{AD}}$ (Hennebelle et al., 2016) for different scenarios, with a fixed MRN dust distribution, accounting for coagulation only, coagulation and electrostatic repulsion, coagulation and erosion, coagulation-repulsion and a lower sticking efficiency ($\epsilon_{\mathrm{stick}}=0.5$). The initial vertical magnetic field is assumed to be $B_{\mathrm{z}}=30\ \mathrm{\mu G}$, and the star-disc system mass to be $M=0.1\ M_{\odot}$. The ambipolar resistivity values $\eta_{\mathrm{AD}}$ are taken at the point of maximum departure from the MRN value, i.e. slightly before $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$.
@@
6.2 Resilient grains or fragile grains: Uncertainties in the collision outcomes
@@
Uncertainties remain in regard to the description of collision outcomes. We recall here that some of them have been ignored in this work (see Wurm & Teiser (2021); Blum (2018) for a detailed list). Bouncing for example, is expected to occur prior to fragmentation, and leads to compaction in place of grain sticking. The bouncing barrier has been widely described in the literature (Zsom et al., 2010). Erosion as a consequence of grain-gas drag could replenish the system with small grains but is expected to be relevant mainly in protoplanetary discs where dust grains can experience a strong gas head wind (Rozner et al., 2020). More relevant to our environment is grain-grain erosion (Gundlach & Blum, 2015; Schrapler et al., 2018) whose influence has been investigated in this work. We based our model on the experimental results of Schrapler et al. (2018). In this regard, we found that grain-grain erosion is unconditionally very efficient at high density, where collisional velocities and size ratios are large, leading to noticeable variations in the Ohm and ambipolar resistivities. However, it is efficient at low density only for sufficiently strong ambipolar drifts. Therefore, given the essential role of this process, future laboratory experiments will have to reinforce our understanding of this collision outcome and to further constrain the velocity threshold and efficiency. Additionally, electrostatic and magnetic reactions resulting from the charges carried by the grains may affect dust growth. In UV-shielded regions (where the grain photo-ionisation is negligible), the electrons are abundant ($n_{\mathrm{e}}>n_{\mathrm{dust}}$, i.e. efficient gas ionisation), plasma charging dominates, and dust grains are expected to be negatively charged, on average (Draine & Sutin, 1987), electrostatic forces thus being repulsive and detrimental to grain coagulation (Okuzumi, 2009; Akimkin et al., 2020). When accounting for dust fragmentation and electrostatic repulsion, the population of sub-micron grains was found to be enhanced and the dust coagulation/fragmentation equilibrium to be severely affected (see Akimkin et al. (2023) who consider hot UV-shielded regions of protoplanetary discs where gas thermal ionisation is important, with $T=1000\ \mathrm{K}$ and $\rho=10^{-12}\ \mathrm{g\ cm^{-3}}$). In the context of protostellar collapse, since the densities involved are very low, we believe that between two consecutive grain-grain collisions, the dust particles would have sufficient time to settle back to their average equilibrium negative charge value imposed by their interaction with electrons and ions. In this case, dust grains would feel repulsive electrostatic forces. As discussed previously, this can curtail to some extent the very efficient coagulation observed at the very beginning of the collapse.
@@
Moreover, the conditions needed for sticking and fragmentation are diverse. Those relevant to a dense core environment are not well constrained in the literature either, and not well reproduced in laboratory experiments. That is why we explored a wide range of grain surface energies $\gamma_{\mathrm{grain}}$. For instance, collision outcomes for icy-grains have turned out to be temperature dependent (Gärtner et al., 2017). Also, icy-grains have been believed to be stickier that bare-silicate grains (Gundlach et al., 2011; Drazkowska, 2017), thus enhancing coagulation and being less prone to bouncing and compaction. Nevertheless, recent studies revealed that the grain surface energy $\gamma_{\mathrm{grain}}$ (which controls both the sticking efficiency and the fragmentation threshold) of icy-grains could sharply decrease from $T=200\ \mathrm{K}$ to $T=175\ \mathrm{K}$ due to a change in the crystal structure (again within the frame of the laboratory experiments, where $P=1\ \mathrm{bar}$ and consequently the ice melting temperature being $T=273\ \mathrm{K}$). Their measurements suggested a grain surface energy ranging between $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$ at low temperature and $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$ at higher temperature, which are the limit values we used in this work. For more details, refer to Musiolik & Wurm (2019). Note however that amorphous ice at low pressure and temperature could recover an efficient sticking under UV irradiation (Musiolik, 2021). We explored the impact of the grain sticking efficiency in Sect. 17 and found that values above $0.3$ do not prevent the population of small grains from being depleted at low densities, thus allowing the ambipolar resistivity $\eta_{\mathrm{AD}}$ to dwell on very large values.